A light beam travels at 1.34×108 m/s in diamond. The wavelength of the light in diamond is 246 nm

Part A

What is the index of refraction of diamond at this wavelength?

Part B

If this same light travels through air, what is its wavelength there?
Answer #1

(a)
Let the refractive index of diamond be, n
v = 1.34 * 10^8 m/s
v = c/n
(1.34 * 10^8) = (3.0*10^8)/n
n = 2.24

(b)
f = c/λ
f = (1.34*10^8)/246 nm
f = 544715.5 * 10^-9 Hz

v = f*λ
λ = v/f
λ = (3.0*10^8)/(544715.5)
λ = 550.75 nm